Circular No. 8463

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 A1 (LINEAR) R. Kracke reports the LINEAR discovery of a comet (discovery ob-

servation below). After posting on the ‘NEO Confirmation Page’, other observers have confirmed the cometary nature on CCD images, including C. Jacques and E. Pimentel (Belo Horizonte, Brazil; 0.30-m reflector, Jan. 14.3 UT; coma diameter ∼ 15′′ and a noticeable 75′′-long tail in p.a. 313^{o}) and J. Young (Table Mountain 0.6-m reflector, Jan. 14.5, moderate cirrus clouds; slightly elongated 12′′ coma of mag 15.0 and a 1′-long broad tail in p.a. 295^{o}).

2005 UT  α_{2000}  δ_{2000}  Mag.
Jan. 13.48135  13^{h}51^{m}08.^{s}43  −17^{o}22′57′′.6  15.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-A56.

T = 2005 Apr. 10.933 TT  ω = 261.^{o}840
Ω = 356.082 }2000.0
q = 1.07894 AU  i = 69.490

SUPERNOVA 2005C T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 18.4) on unfiltered CCD images taken on Jan. 11.051 and 13.036 UT. The new object is located at α = 11^{h}15^{m}29^{s}.42, δ = +60^{o}45′05′′.5 (equinox 2000.0), which is ≈ 3′′.0 west and 10′′.7 north of the center of its apparent host galaxy. SN 2005C is not present on Boles’ images from 2004 Apr. 21 and Nov. 20 (limiting mag 19.5) or on Digitized Sky Survey plates from 1993 (limiting red mag 21.0) or 1994 (limiting blue mag 21.5).

SUPERNOVA 2004gy Further to IAUC 8461, M. Modjaz, R. Kirshner, and P. Challis report

that a spectrogram of SN 2004gy (cf. IAUC 8460), obtained at Mt. Hopkins by W. Brown on Jan. 10.49 UT, reveals it to be a type-II supernova, probably within a few weeks past explosion. The spectrum consists of a blue continuum and Balmer-line P-Cyg features, with an emission-dominated Hα component. The host-galaxy redshift, measured from narrow Hα emission lines in the nucleus, is 0.0269; adopting this redshift, the expansion velocity derived from the minimum of the Hβ line is ∼ 10000 km/s.

2005 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green